Indicate whether each structure is aromatic, nonar
Answer #1

BH O aromatic O nonaromatic! O nonaromatic! |○ nonaromatic | |O nonaromatic| |O nonaromatic aromatic aromatic aromatic O arom

Explanation

1) For compound A, number of \pi electrons = 6. Therefore 4n+2 rule and aromatic

2) For compound B, number of \pi electrons = 6. Therefore 4n+2 rule and aromatic

3) For compound C, number of \pi electrons = 8. Therefore 4n rule and anti-aromatic

4) For compound D, number of \pi electrons = 6. Therefore 4n+2 rule and aromatic

5) For compound E, number of \pi electrons = 8. Therefore 4n rule and anti-aromatic